# TelePreview: A User-Friendly Teleoperation System with Virtual Arm Assistance for Enhanced Effectiveness

### v4

@@ VI-F Deployment on Various Robots (Q3)
We deploy our TelePreview system on a Ufactory xArm [40] equipped with both a LEAP hand [34] and a parallel gripper. By updating only the kinematic parameters and transformation chain for each end-effector, TelePreview preserves its full functionality across diverse setups, as shown in Fig. 8. This successful deployment on multiple end-effector configurations illustrates the system’s adaptability to new hardware. Its modular design allows users to seamlessly integrate TelePreview into different robotic platforms while retaining all key features—preview, multi-view visualization, and intuitive control interfaces.
@@ VII Conclusion
In this paper, we presented TelePreview, a low-cost teleoperation system featuring AR preview feedback designed for deployment across diverse robotic platforms. Our experiments demonstrate that TelePreview offers sufficient flexibility and safety to successfully perform a variety of fine-grained tasks. TelePreview provides immediate, robot-centric feedback that enables users to preview and refine their commands before physical execution.
While our results show promising advances in teleoperation interfaces, a key limitation is the visual ambiguity caused by occlusions between the preview robot and scene objects. Future work could address this by incorporating depth information from RGB-D cameras to create more accurate spatial relationships between virtual and physical elements.

### v6

@@ VI-F Deployment on Various Robots (Q3)
We evaluate the generalizability of TelePreview across three end-effector types on a Ufactory xArm: a 16-DoF anthropomorphic LEAP hand, a parallel-jaw gripper, and a vacuum gripper (both 1-DoF). Our results (see Appendix) show that TelePreview yields the most significant performance gains with the LEAP hand, highlighting its effectiveness for high-DoF, fine-grained manipulation tasks. Improvements with the gripper and vacuum gripper are more modest but still demonstrate the system’s utility across diverse hardware configurations. These findings support our claim that TelePreview generalizes well while providing the greatest benefit in dexterous settings.
@@ VII Conclusion
In this paper, we presented TelePreview, a low-cost teleoperation system featuring AR preview feedback designed for deployment across diverse robotic platforms. Our experiments demonstrate that TelePreview offers sufficient flexibility and safety to perform a variety of fine-grained tasks successfully. TelePreview provides immediate, robot-centric feedback that enables users to preview and refine their commands before physical execution.
While our results show promising advances in teleoperation interfaces, a key limitation is the visual ambiguity caused by occlusions between the preview robot and scene objects. This limitation stems from our use of alpha blending, which lacks depth-aware occlusion reasoning and cannot distinguish whether the virtual arm should occlude real objects or vice versa, resulting in artifacts even on passthrough-capable devices like the Meta Quest 3 without real-time depth estimation. To mitigate this ambiguity, our approach employs multi-view overlays (e.g., top-down and side views), allowing operators to compare perspectives and correctly infer spatial ordering.
Future work will explore incorporating depth estimation from RGB-D or passthrough feeds and using it to inform depth-aware rendering and spatial reasoning. These could reduce user uncertainty during manipulation, leading to cleaner demonstration data and a more intuitive teleoperation experience.